A board 65 feet long is cut into two pieces such that the length of one piece is two-thirds of the length of the other piece. Find the length of the shorter piece of board.
Answer by stanbon(75887)

A board 65 feet long is cut into two pieces such that the length of one piece is two-thirds of the length of the other piece. Find the length of the shorter piece of board.
------
Equations:
x + y = 65 feet
x = (2/3)y
=====
Substitute and solve for "y":
(2/3)y + y = 65
(5/3)y = 65
(1/3)y = 13
y = 39 feet
=====
Solve for "x":
x = (2/3)y
x = (2/3)39
x = 26 feet
==================
Cheers,
Stan H.
================
x =